Skip to main contentSkip to solution

Suppose X has Poisson distribution such that 3P(X=1)=2P(X=2)3 P(X=1) = 2 P(X=2) then P(X>0)P(X>0) is:

Solution

✅ Correct Option: 3

For a Poisson distribution with parameter λ:

P(X=k)=e−λ⋅λkk!P(X = k) = \frac{e^{-\lambda} \cdot \lambda^k}{k!}

Given condition: 3P(X=1)=2P(X=2)3P(X=1) = 2P(X=2)


For X=1X = 1:

P(X=1)=e−λ⋅λ11!P(X=1) = \frac{e^{-\lambda} \cdot \lambda^1}{1!}

P(X=1)=λe−λP(X=1) = \lambda e^{-\lambda}

For X=2X = 2:

P(X=2)=e−λ⋅λ22!P(X=2) = \frac{e^{-\lambda} \cdot \lambda^2}{2!}

P(X=2)=λ2e−λ2P(X=2) = \frac{\lambda^2 e^{-\lambda}}{2}


Using the given condition:

3(λe−λ)=2(λ2e−λ2)3(\lambda e^{-\lambda}) = 2\left(\frac{\lambda^2 e^{-\lambda}}{2}\right)

3λe−λ=λ2e−λ3\lambda e^{-\lambda} = \lambda^2 e^{-\lambda}

Dividing both sides by λe−λ\lambda e^{-\lambda}:

3=λ3 = \lambda


Using the complement rule:

P(X>0)=1−P(X=0)P(X>0) = 1 - P(X=0)

For λ=3\lambda = 3:

P(X=0)=e−3⋅300!P(X=0) = \frac{e^{-3} \cdot 3^0}{0!}

P(X=0)=e−3⋅11P(X=0) = \frac{e^{-3} \cdot 1}{1}

P(X=0)=e−3P(X=0) = e^{-3}


Therefore:

P(X>0)=1−e−3P(X>0) = 1 - e^{-3}

Keyboard Shortcuts

  • Left arrow: Previous question
  • Right arrow: Next question
  • S key: Jump to solution
  • Q key: Jump to question